### v2

# Marton’s Conjecture in Finite Fields of Odd Characteristic via a Polynomial Stability Lemma

Abstract
We settle the odd-characteristic finite-field case of Marton’s conjecture by proving a sharp polynomial stability dichotomy: either the $L^{4}$ Fourier mass of $\mathbf{1}_{A}$ concentrates on a span of dimension $\mathrm{poly}(K)$, or in a quotient of codimension $\mathrm{poly}(K)$ the doubling constant decreases by at least $K^{-C}$. As a consequence we obtain a polynomial Freiman–Ruzsa (PFR/Marton) covering theorem in $\mathbb{F}_{p}^{n}$ for $p$ odd. Together with the characteristic $2$ result of Green, Gowers, Manners, and Tao [1], this yields a complete resolution of the finite-field formulation of Marton’s conjecture across all characteristics.
1 Introduction
Let $G$ be an abelian group and $A\subset G$. Marton’s conjecture (also known as the polynomial Freiman-Ruzsa conjecture) asserts that small doubling, $|A+A|\leq K|A|$, forces $A$ to be covered by $\mathrm{poly}(K)$ cosets of a subgroup whose index also depends polynomially on $K$. This conjecture has been a central problem in additive combinatorics, with connections to sum-product phenomena, incidence geometry, and arithmetic progressions.
1.1 Context and related work
Recently, Green, Gowers, Manners, and Tao [1] resolved the case of characteristic $2$, establishing the conjecture for subsets of $\mathbb{F}_{2}^{n}$ with polynomial bounds in the doubling constant $K$. Their proof introduced a novel entropy-increment strategy combined with structure theorems for sets with small doubling in characteristic $2$. However, the techniques are fundamentally tied to the characteristic $2$ setting and do not directly extend to odd characteristic, where different phenomena arise in the Fourier-analytic landscape.
Concurrent and subsequent work.
We note that, subsequent to the initial development of this work, Gowers, Green, Manners, and Tao [2] established Marton’s conjecture for abelian groups of bounded torsion using entropy-theoretic methods. Since finite fields are groups of bounded torsion, their result also covers finite fields of odd characteristic. The present work provides an independent, direct Fourier-analytic proof with explicit spectral stability dichotomies and quantitative polynomial bounds in the finite-field setting.
The characteristic $2$ case exhibits special structural features: notably, the additive energy can be analyzed via quadratic Fourier analysis, and certain cancellations in the convolution algebra simplify the entropy arguments. In contrast, odd characteristic requires different methods to control the spectral dispersion and quotient improvements.
In particular, in odd characteristic the lack of characteristic-$2$ cancellations means that an entropy increment alone does not force a single large spectrum to dominate; one must instead track how $L^{4}$ Fourier mass disperses and show that dispersion creates a genuine quotient improvement. This is the conceptual role of the polynomial stability lemma.
1.2 Our contribution
We resolve the odd characteristic case by developing a polynomial stability lemma (PSL) based on spectral concentration and quotient improvement. Our approach uses:
  • Chang-type dissociation bounds to control the dimension of large-spectrum spans (Lemma 4.1);
  • Paley-Zygmund concentration to convert spectral mass into near-coset structure (Lemma 5.1);
  • Two-scale Fourier projection to extract polynomial quotient improvements when spectral mass is dispersed (Lemma 6.4).
Combined with the result of [1], this completes the resolution of Marton’s conjecture for all finite fields.
1.3 Main theorem
Theorem 1.1.
There exists an absolute constant $C\geq 1$ such that for any nonempty $A\subset\mathbb{F}_{p}^{n}$ ($p$ an odd prime) with $|A+A|\leq K|A|$ there exist a subgroup $H\leq\mathbb{F}_{p}^{n}$ and translates $x_{1}+H,\dots,x_{m}+H$ with
$m\leq K^{C},\qquad|H|\leq K^{C}|A|,\qquad A\subset\bigcup_{j=1}^{m}(x_{j}+H).$
Our proof proceeds via a polynomial stability lemma (PSL, Lemma 2.1) that provides a dichotomy: either $A$ is almost contained in a single coset of a subgroup of index $\mathrm{poly}(K)$, or in a quotient of codimension $\mathrm{poly}(K)$ the doubling constant decreases by at least $K^{-C}$. Iterating the second alternative yields the first in $\mathrm{poly}(K)$ steps, proving Theorem 1.1.
1.4 Proof roadmap
After setting Fourier-analytic notation, we proceed in three stages:
  1. (i)
    We fix a spectral threshold $\tau=K^{-c_{0}}$ and let $V$ be the span of the large spectrum. Lemma 4.1 shows $\dim V\ll\mathrm{poly}(K)$.
  2. (ii)
    We decompose $\mathbf{1}_{A}=f_{\mathrm{low}}+f_{\mathrm{high}}$ via projection to $V$. If the $L^{4}$ mass concentrates on $V$, Lemma 5.1 upgrades this to a near-coset inside $H=V^{\perp}$.
  3. (iii)
    If the $L^{4}$ mass is dispersed, a two-scale dissociated extraction in $\widehat{G}/V$ produces a complementary span $V^{\prime}$ of dimension $\mathrm{poly}(K)$. In the quotient by $H^{\prime}=(V^{\prime})^{\perp}$ we detect a large Fourier coefficient, convert it to an energy boost, and hence to a polynomial decrement of the doubling constant (Lemma 6.4). Iterating this improvement yields PSL and thus Theorem 1.1.
1.5 Notation
Fourier conventions (normalization).
We fix $\widehat{f}(\xi)=|G|^{-1}\sum_{x\in G}f(x)e_{p}(-\langle x,\xi\rangle)$ and $(f*g)(x)=\sum_{y\in G}f(y)g(x-y)$. With this choice, Parseval is $\|f\|_{2}^{2}=|G|\sum_{\xi\in\widehat{G}}|\widehat{f}(\xi)|^{2}$ and $\mathsf{E}(A)=|G|\sum_{\xi\in\widehat{G}}|\widehat{\mathbf{1}_{A}}(\xi)|^{4}$ as in (1). For a quotient $\pi:G\to G/H^{\prime}$ with $(H^{\prime})^{\perp}=V^{\prime}$, every nontrivial $\chi$ on $G/H^{\prime}$ corresponds to a unique $\xi\in V^{\prime}$; then $\widehat{\mathbf{1}_{\pi(A)}}(\chi)=\widehat{\mathbf{1}_{A}}(\xi)$. We use this in Claim 6.2.
We write $\alpha:=|A|/|G|$ and denote by $\mathbf{1}_{A}$ the indicator function of $A$.
Expectations and variance.
We write $\mathbb{E}[\cdot]$ for expectation with respect to the uniform measure on $G$ (or on the relevant quotient), so $\mathbb{E}[g]=\alpha$ and $\operatorname{Var}(g)=\mathbb{E}[g^{2}]-(\mathbb{E}[g])^{2}$. The normalized Fourier transform on an abelian group $G$ is
$\widehat{f}(\xi):=\frac{1}{|G|}\sum_{x\in G}f(x)\,e_{p}(-\langle x,\xi\rangle),\qquad e_{p}(t):=e^{2\pi it/p}.$
For $A\subset G$, the additive energy is
$\mathsf{E}(A)=\sum_{x}r_{A+A}(x)^{2}=\left\lVert\mathbf{1}_{A}*\mathbf{1}_{A}\right\rVert_{2}^{2}=|G|\sum_{\xi\in\widehat{G}}|\widehat{\mathbf{1}_{A}}(\xi)|^{4}.$ (1)
By Cauchy-Schwarz, small doubling $|A+A|\leq K|A|$ implies the lower bound
$\mathsf{E}(A)\ \geq\ \frac{|A|^{4}}{|A+A|}\ =\ \frac{|A|^{3}}{K}.$ (2)
2 Polynomial Stability Lemma (PSL)
Convention on constants.
We use $C,C_{1},C_{2},\dots$ for absolute positive constants whose values may change from line to line, and $c,c_{0},c_{1},\dots$ for absolute positive constants that may be taken sufficiently small. Dependencies are indicated explicitly when needed (e.g. $C=C(c_{0},c)$). Parameters depending on the doubling constant $K$ are named (e.g. $\tau=K^{-c_{0}}$, $\beta=K^{-c}$), while the final exponents in all theorems and lemmas are independent of $K$. Throughout, $\dim V\ll K^{C_{1}}$ abbreviates $\dim V\leq C^{\prime}K^{C_{1}}$ for an absolute $C^{\prime}>0$.
The following dichotomy is the core of our argument.
Lemma 2.1 (Polynomial Stability Lemma).
Fix $\varepsilon\in(0,1/10)$. There exists $C=C(\varepsilon)\geq 1$ such that for any nonempty $A\subset\mathbb{F}_{p}^{n}$ with $|A+A|\leq K|A|$ and $K\geq 2$, one of the following holds:
  • (1)
    Near-coset structure: There exists a subgroup $H\leq\mathbb{F}_{p}^{n}$ and a coset $x+H$ with
    $|H|\leq K^{C}|A|\quad\text{and}\quad|A\cap(x+H)|\geq(1-\varepsilon)|H|.$
  • (2)
    Polynomial improvement in a quotient: There exists a subgroup $H^{\prime}\leq\mathbb{F}_{p}^{n}$ with $\operatorname{codim}(H^{\prime})\leq K^{C}$ such that, writing $\pi:\mathbb{F}_{p}^{n}\to\mathbb{F}_{p}^{n}/H^{\prime}$ and $A^{\prime}=\pi(A)$, we have
    $|A^{\prime}+A^{\prime}|\leq(K-\delta)|A^{\prime}|,\quad|A^{\prime}|\geq K^{-C}|A|,\quad\text{and}\quad\delta\geq K^{-C}.$
Proof that PSL implies Theorem 1.1.
Suppose we are in case (2). Then in the quotient $\mathbb{F}_{p}^{n}/H^{\prime}$, the set $A^{\prime}$ has doubling constant at most $K-\delta\geq K-K^{-C}$. Iterating this process, after at most $O(K/\delta)\leq O(K^{C+1})$ steps the doubling constant drops below an absolute threshold, forcing case (1). The total codimension accumulated is at most $K^{C}\cdot K^{C+1}\leq K^{C^{\prime}}$ for some $C^{\prime}=C^{\prime}(\varepsilon)$. Once we reach case (1), standard covering arguments (partitioning the complement and applying Ruzsa calculus) yield the conclusion of Theorem 1.1. ∎
The remainder of this paper is devoted to proving Lemma 2.1.
A monotone invariant in the main text
For completeness we record a self-contained version of the iteration potential used to deduce Theorem 1.1 from PSL.
Proposition 2.2 (Monotone potential in the main text).
Let $(A_{j}\subset G_{j})$ be obtained by iterating the improvement alternative of Lemma 2.1, with $K_{j}:=|A_{j}+A_{j}|/|A_{j}|$ and $\alpha_{j}:=|A_{j}|/|G_{j}|$. Then there exists an absolute $\gamma\geq 1$ such that
$\mathcal{I}_{j}:=K_{j}\,\alpha_{j}^{-\gamma}$
is nonincreasing at each improvement step. Consequently, the number of improvement steps is $O(K_{0}^{C+1})$ and the total accumulated codimension is $K_{0}^{C^{\prime}}$ for some absolute $C^{\prime}>0$.
Proof.
In an improvement step we have $K_{j+1}\leq K_{j}-\delta_{j}$ with $\delta_{j}\geq K_{j}^{-C}$ and $|A_{j+1}|\geq K_{j}^{-C}|A_{j}|$, so $\alpha_{j+1}\geq K_{j}^{-C}\alpha_{j}$. Hence
$\mathcal{I}_{j+1}\ \leq\ (K_{j}-\delta_{j})\,(K_{j}^{-C}\alpha_{j})^{-\gamma}\ =\ \mathcal{I}_{j}\cdot K_{j}^{C\gamma}\Big(1-\frac{\delta_{j}}{K_{j}}\Big).$
Choosing $\gamma$ (e.g. $\gamma=C+2$) ensures $K_{j}^{C\gamma}(1-\delta_{j}/K_{j})\leq 1$ for all $K_{j}\geq 2$, giving monotonicity. Summing decrements gives the step bound; codimension accumulates to at most $K_{0}^{C^{\prime}}$ since each step adds at most $K_{0}^{C}$. ∎
3 Spectral setup and two-scale projection
Lemma 3.1 (Dissociated extraction in $\widehat{G}/V$).
Let $V\leq\widehat{G}$ and let $S\subset V^{c}$. Let $\overline{\widehat{G}}:=\widehat{G}/V$ and write $\overline{S}=\{\overline{\xi}:\xi\in S\}\subset\overline{\widehat{G}}$. If $\overline{\Xi}\subset\overline{S}$ is a maximal dissociated set in $\overline{\widehat{G}}$ and $\Xi$ are chosen representatives in $S$, then $V^{\prime}:=\operatorname{Span}(\Xi)$ satisfies $V^{\prime}\cap V=\{0\}$ and $\dim V^{\prime}=\dim\operatorname{Span}(\overline{\Xi})$. Moreover, if $\overline{S}$ obeys a Chang-type bound $|\overline{\Xi}|\ll\tau^{-2}\log(1/\alpha)$, then $\dim V^{\prime}\ll\tau^{-2}\log(1/\alpha)$.
Proof.
If a nontrivial linear combination of $\Xi$ landed in $V$, then the corresponding combination of classes would vanish in $\overline{\widehat{G}}$, contradicting dissociation. The dimension and quantitative bound follow immediately. ∎
For $\tau\in(0,1)$, define the large spectrum
$\operatorname{Spec}_{\tau}(A):=\{\xi\in\widehat{G}:|\widehat{\mathbf{1}_{A}}(\xi)|\geq\tau\alpha\}.$
Let $V:=\operatorname{Span}(\operatorname{Spec}_{\tau}(A))\leq\widehat{G}$ be the span of the large-spectrum characters, and let $H:=V^{\perp}\leq\mathbb{F}_{p}^{n}$ be its annihilator. Let $P$ denote convolution with the uniform measure $\mu_{H}$ on $H$ (this is the exact projector onto $V$ in Fourier space) and let Q=IPQ=I-P. We decompose (with $f_{\mathrm{low}}$ the $V$-projected component and $f_{\mathrm{high}}$ orthogonal to $V$)
$\mathbf{1}_{A}\ =\ f_{\mathrm{low}}+f_{\mathrm{high}},\qquad f_{\mathrm{low}}:=P\mathbf{1}_{A},\quad f_{\mathrm{high}}:=Q\mathbf{1}_{A}.$
In Fourier space, $\widehat{f}_{\mathrm{low}}=1_{V}\cdot\widehat{\mathbf{1}_{A}}$ and $\widehat{f}_{\mathrm{high}}=1_{V^{c}}\cdot\widehat{\mathbf{1}_{A}}$.
4 Lemma S1: Polynomial Chang span bound
Lemma 4.1.
There exist absolute constants $c_{0},C_{1}>0$ such that, for $\tau:=K^{-c_{0}}$, we have
$\dim V\ \ll\ K^{C_{1}}.$
Moreover, one may take $C_{1}\leq C_{\mathrm{RC}}\cdot 2c_{0}$ where $C_{\mathrm{RC}}$ is an absolute constant from the Rudin–Chang inequality.
Proof.
Let $D\subset\operatorname{Spec}_{\tau}(A)$ be a maximal dissociated set. By the Rudin-Chang inequality (see [3, 4]), there exists an absolute constant $c>0$ such that
$\sum_{\eta\in D}|\widehat{\mathbf{1}_{A}}(\eta)|^{2}\ \leq\ c\,\alpha^{2}\log\frac{1}{\alpha}.$
Since each $\eta\in D$ lies in $\operatorname{Spec}_{\tau}(A)$, we have $|\widehat{\mathbf{1}_{A}}(\eta)|\geq\tau\alpha$. Therefore,
$|D|\cdot\tau^{2}\alpha^{2}\ \leq\ \sum_{\eta\in D}|\widehat{\mathbf{1}_{A}}(\eta)|^{2}\ \leq\ c\,\alpha^{2}\log\frac{1}{\alpha},$
which yields
$|D|\ \leq\ c\,\tau^{-2}\log\frac{1}{\alpha}.$
By the standing assumptions (Appendix E) and the modeling Lemma A.1, we may, without loss of generality, work in a Freiman model where the density satisfies $\alpha\in[K^{-C},1/2]$. In particular $\log(1/\alpha)\ll\log K$. Choosing $\tau=K^{-c_{0}}$ for a suitably small $c_{0}>0$, we obtain
$|D|\ \ll\ K^{2c_{0}}\log K\ \ll\ K^{C_{1}}$
for some absolute $C_{1}>0$. Since $V$ is spanned by the dissociated set $D$, we have $\dim V\leq|D|\ll K^{C_{1}}$. ∎
5 Lemma S3: From spectral mass to a near-coset
Lemma 5.1.
Suppose that for some $c\in(0,1)$ we have
$\sum_{\xi\in V}|\widehat{\mathbf{1}_{A}}(\xi)|^{4}\ \geq\ c\sum_{\xi\in\widehat{G}}|\widehat{\mathbf{1}_{A}}(\xi)|^{4}.$
Then there exists a coset $x+H$ with
$|A\cap(x+H)|\ \geq\ (1-\varepsilon)|H|,$
provided $\varepsilon\leq\tfrac{1}{2}\sqrt{cc_{1}}$ where $c_{1}>0$ is the implied constant in (3).
Proof.
Parseval gives
$\left\lVert\mathbf{1}_{A}*\mu_{H}\right\rVert_{2}^{2}\ =\ |G|\sum_{\xi\in V}|\widehat{\mathbf{1}_{A}}(\xi)|^{2}\ \geq\ c_{1}\,\alpha^{2}|G|$ (3)
for an absolute $c_{1}=c_{1}(c)>0$ by Hölder together with the hypothesis and (1). Let $g:=\mathbf{1}_{A}*\mu_{H}$, so $g(x)\in[0,1]$ and $\mathbb{E}[g]=\alpha$. Then $\mathbb{E}[g^{2}]\geq(1+c_{1})\alpha^{2}$ and $\operatorname{Var}(g)=\mathbb{E}[g^{2}]-(\mathbb{E}[g])^{2}\geq c_{1}\alpha^{2}$. By Paley–Zygmund,
$\mathbb{P}\{g\geq(1-\varepsilon)\alpha\}\ \geq\ \frac{(\varepsilon\alpha)^{2}}{\mathbb{E}[g^{2}]}\ \geq\ \frac{\varepsilon^{2}}{1+c_{1}}.$
Choosing $\varepsilon\leq\tfrac{1}{2}\sqrt{cc_{1}}$ ensures a positive lower bound, yielding a coset with $g(x)\geq(1-\varepsilon)$. ∎
6 Lemma S2: Dispersion and polynomial improvement
Lemma 6.1 (Energy-to-doubling transfer without loss of scale).
Let $A^{\prime}\subset G^{\prime}$ with density $\alpha^{\prime}$, and suppose there exists a nontrivial character $\chi$ on $G^{\prime}$ with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\geq\eta\,\alpha^{\prime}$ for some $\eta\in(0,1]$. Then
$\mathsf{E}(A^{\prime})\ \geq\ \alpha^{\prime 4}|G^{\prime}|\,(1+c_{*}\eta^{4})$ (4)
for an absolute constant $c_{*}>0$. Consequently, if $|A^{\prime}+A^{\prime}|\leq K\,|A^{\prime}|$ then in fact $|A^{\prime}+A^{\prime}|\leq(K-\delta)|A^{\prime}|$ with $\delta\geq c_{*}\eta^{4}$.
Proof.
Plancherel gives $\mathsf{E}(A^{\prime})=|G^{\prime}|\sum_{\psi}|\widehat{\mathbf{1}_{A^{\prime}}}(\psi)|^{4}$. The trivial character contributes $\alpha^{\prime 4}|G^{\prime}|$. If there is a nontrivial $\chi$ with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\geq\eta\alpha^{\prime}$, the sum gains at least $\eta^{4}\alpha^{\prime 4}$ from that term alone, hence (4) with $c_{*}=1$. The inequality $|A^{\prime}|^{4}\leq\mathsf{E}(A^{\prime})\,|A^{\prime}+A^{\prime}|$ then yields the claimed decrement. ∎
We first establish a technical claim controlling Fourier coefficients under quotient maps when projections are mismatched.
Claim 6.2 (Exact Fourier control under quotient via dissociation modulo $V$).
Let $P$ be the projector onto $V\leq\widehat{G}$ (convolution by $\mu_{H}$ with $H=V^{\perp}$) and set $f=\mathbf{1}_{A}$, B=(IP)fB=(I-P)f. Let $\overline{\widehat{G}}:=\widehat{G}/V$ be the quotient dual group and write $\overline{\xi}$ for the class of $\xi\in\widehat{G}$ modulo $V$. Let $S:=\{\xi\in V^{c}:|\widehat{B}(\xi)|\geq\lambda\}$ for some threshold $\lambda>0$, and let $\overline{\Xi}\subset\overline{S}:=\{\overline{\xi}:\xi\in S\}\subset\overline{\widehat{G}}$ be a maximal dissociated set in the quotient. Pick representatives $\Xi=\{\xi_{1},\dots,\xi_{r}\}\subset S$ of $\overline{\Xi}$ and set $V^{\prime}:=\operatorname{Span}(\Xi)$ and $H^{\prime}=(V^{\prime})^{\perp}$.
Then $V^{\prime}\cap V=\{0\}$ and for every nontrivial character $\chi$ on $G/H^{\prime}$ (i.e., $\chi\leftrightarrow\xi\in V^{\prime}\setminus\{0\}$) we have exactly
$\widehat{f^{\prime}}(\chi)\ =\ \widehat{B^{\prime}}(\chi)\ =\ \widehat{B}(\xi),$
where $f^{\prime}=\mathbf{1}_{A^{\prime}}$ for $A^{\prime}=\pi(A)$ with $\pi:G\to G/H^{\prime}$ and $B^{\prime}$ is defined by $B^{\prime}\circ\pi=B*\mu_{H^{\prime}}$. In particular, for all such $\chi$ we also have $|\widehat{f^{\prime}}(\chi)|=|\widehat{B}(\xi)|$.
Remark 6.3 (Uniqueness of lift and choice of representatives).
Working modulo $V$ in $\widehat{G}/V$ ensures $V^{\prime}\cap V=\{0\}$, so each nontrivial $\chi$ on $G/H^{\prime}$ lifts to a unique $\xi\in V^{\prime}$ and $\widehat{f^{\prime}}(\chi)=\widehat{f}(\xi)$ is well-defined. We choose representatives $\Xi$ of a maximal dissociated subset of $\overline{S}=\{\overline{\xi}\in\widehat{G}/V:\ |\widehat{B}(\xi)|\geq\lambda\}$ so that $|\widehat{f^{\prime}}(\chi)|=|\widehat{B}(\xi)|\geq\lambda$ for all nontrivial $\chi$ arising from $V^{\prime}$. This removes aliasing between $V$ and $V^{\prime}$.
Proof.
By construction $\overline{\Xi}$ is dissociated in $\overline{\widehat{G}}$, so $\operatorname{Span}(\overline{\Xi})\cap\{\overline{0}\}=\{\overline{0}\}$. Equivalently, $\operatorname{Span}(\Xi)\cap V=\{0\}$, hence $V^{\prime}\cap V=\{0\}$. Therefore $\widehat{\mu_{H}}(\xi)=0$ for every nonzero $\xi\in V^{\prime}$ and so $\widehat{B}(\xi)=\widehat{f}(\xi)$ on $V^{\prime}\setminus\{0\}$. Since $\widehat{\mu_{H^{\prime}}}(\xi)=1$ for $\xi\in V^{\prime}$, we have $\widehat{B^{\prime}}(\chi)=\widehat{B}(\xi)$ and $\widehat{f^{\prime}}(\chi)=\widehat{f}(\xi)$ where $\chi$ is the character on $G/H^{\prime}$ corresponding to $\xi$. Combining these equalities gives $\widehat{f^{\prime}}(\chi)=\widehat{B^{\prime}}(\chi)=\widehat{B}(\xi)$, which is exact. ∎
Lemma 6.4 (S2: Dispersion and polynomial improvement).
There exist absolute constants $c,C>0$ and a choice $\tau=K^{-c}$ with the following dichotomy. Let $V=\operatorname{Span}(\operatorname{Spec}_{\tau}(A))$ and $H=V^{\perp}$.
  • (i)
    (Concentration) $\sum_{\xi\in V}|\widehat{\mathbf{1}_{A}}(\xi)|^{4}\geq(1-K^{-c})\sum_{\xi\in\widehat{G}}|\widehat{\mathbf{1}_{A}}(\xi)|^{4}$.
  • (ii)
    (Improvement) There exists a subspace $V^{\prime}\leq\widehat{G}$ with $V^{\prime}\cap V=\{0\}$ and $\dim V^{\prime}\ll K^{C}$, and $H^{\prime}=(V^{\prime})^{\perp}$, such that for $A^{\prime}=\pi(A)\subset G/H^{\prime}$ we have
    $|A^{\prime}+A^{\prime}|\ \leq\ (K-\delta)|A^{\prime}|,\qquad|A^{\prime}|\ \geq\ K^{-C}|A|,\qquad\delta\ \geq\ K^{-C}.$
Proof.
Assume (i) fails, so the $L^{4}$-mass outside $V$ is at least a $K^{-c}$ fraction of the total. Let $B=(I-P)\mathbf{1}_{A}$ and set the threshold $\lambda:=\tfrac{1}{2}\tau\alpha$. Define $S=\{\xi\in V^{c}:|\widehat{B}(\xi)|\geq\lambda\}$; by the $L^{4}$-mass lower bound and Markov, $S\neq\emptyset$. Extract $\overline{\Xi}\subset\overline{S}\subset\widehat{G}/V$ maximal dissociated and pick representatives $\Xi\subset S$; let $V^{\prime}=\operatorname{Span}(\Xi)$ and $H^{\prime}=(V^{\prime})^{\perp}$. By Lemma 3.1, $V^{\prime}\cap V=\{0\}$ and $\dim V^{\prime}\ll\tau^{-2}\log(1/\alpha)\ll K^{C}$.
Passing to the quotient $G/H^{\prime}$ and writing $A^{\prime}=\pi(A)$, Claim 6.2 yields exact equality $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|=|\widehat{B}(\xi)|$ for every nontrivial $\chi$ coming from $\xi\in V^{\prime}$, hence in particular there exists $\chi_{*}\neq 1$ with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi_{*})|\geq\lambda=\tfrac{1}{2}\tau\alpha$. As $|A^{\prime}|/|G/H^{\prime}|=\alpha^{\prime}\asymp\alpha$ up to $K^{\pm C}$, we have $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi_{*})|\geq K^{-C}\alpha^{\prime}$. Applying Lemma 6.1 with $\eta:=K^{-C}$ gives an additive energy boost of size $c_{*}K^{-4C}\alpha^{\prime 4}|G/H^{\prime}|$, which in turn implies $|A^{\prime}+A^{\prime}|\leq(K-\delta)|A^{\prime}|$ with $\delta\gg K^{-C}$ after adjusting $C$. Finally, from $\dim V^{\prime}\ll K^{C}$ we get $|H^{\prime}|\leq K^{C^{\prime}}$ for some absolute $C^{\prime}$, hence $|A^{\prime}|\geq K^{-C}|A|$. ∎
7 Proof of the Polynomial Stability Lemma
Proof of Lemma 2.1.
By Lemma 4.1, we have $\dim V\ll K^{C_{1}}$ for $\tau=K^{-c_{0}}$. By Lemma 6.4, either we are in the concentration case (i) or we obtain the improvement case (ii). If we are in case (i), then by Lemma 5.1 (choosing $c=1-K^{-c}$ and adjusting $\varepsilon$), there exists a coset $x+H$ with $|A\cap(x+H)|\geq(1-\varepsilon)|H|$ and $|H|\leq K^{C}|A|$ (using $\dim V\ll K^{C_{1}}$), which is PSL case (1). If we are in case (ii), we obtain PSL case (2) directly from Lemma 6.4. ∎
8 Resolution of Marton’s conjecture across all characteristics
Our main result (Theorem 1.1) establishes the polynomial Freiman-Ruzsa conjecture (Marton’s conjecture) for finite fields of odd characteristic. The characteristic $2$ case was resolved by Green, Gowers, Manners, and Tao [1], who proved that for $A\subset\mathbb{F}_{2}^{n}$ with $|A+A|\leq K|A|$, there exists a subgroup $H$ and translates covering $A$ with $m\leq\mathrm{poly}(K)$ and $|H|\leq\mathrm{poly}(K)|A|$.
Combining these results, we obtain:
Theorem 8.1 (Marton’s conjecture for all finite fields).
There exists an absolute constant $C\geq 1$ such that the following holds. Let $\mathbb{F}_{q}$ be any finite field and let $A\subset\mathbb{F}_{q}^{n}$ satisfy $|A+A|\leq K|A|$. Then there exist a subgroup $H\leq\mathbb{F}_{q}^{n}$ and translates $x_{1}+H,\dots,x_{m}+H$ with
$m\leq K^{C},\qquad|H|\leq K^{C}|A|,\qquad A\subset\bigcup_{j=1}^{m}(x_{j}+H).$
Proof.
If $\mathrm{char}(\mathbb{F}_{q})=2$, this is the main result of [1]. If $\mathrm{char}(\mathbb{F}_{q})$ is odd, this is Theorem 1.1 of the present paper. ∎
Remark 8.2.
The constant $C$ in Theorem 8.1 may differ between characteristic $2$ and odd characteristic, but both are absolute (independent of $q$, $n$, and $K$). The techniques used in the two cases are fundamentally different: the characteristic $2$ proof of [1] relies on entropy-increment methods, while our odd-characteristic proof uses spectral concentration and quotient improvements.
Remark 8.3.
The question of whether a universal polynomial bound (with the same exponent $C$ for all characteristics) exists remains open. However, for applications, the existence of characteristic-dependent polynomial bounds suffices.
9 Concluding remarks
9.1 Comparison of methods
The characteristic $2$ approach of [1] and our odd-characteristic method exhibit interesting parallels and contrasts:
  • Entropy vs. energy: The characteristic $2$ proof uses an entropy-increment argument to force structure, exploiting special properties of quadratic Fourier analysis in $\mathbb{F}_{2}^{n}$. Our approach uses $L^{4}$-Fourier energy concentration and dispersion, which is more natural in the odd-characteristic setting where higher-order Fourier analysis is less constrained.
  • Stability dichotomies: Both proofs proceed via dichotomies (concentration vs. improvement), but the mechanisms differ. In characteristic $2$, the dichotomy arises from entropy comparisons; in odd characteristic, it arises from spectral mass distribution.
  • Polynomial bounds: Both methods yield polynomial bounds in $K$, but the exponents may differ. Our constants $c,C$ are absolute and can be made explicit by tracking the dissociation bounds and Fourier estimates.
9.2 Further directions
Several natural questions arise from our work:
  1. 1.
    Optimal exponents: What is the best polynomial bound achievable? Can the exponent $C$ in Theorem 1.1 be improved, and is there a fundamental lower bound?
  2. 2.
    General abelian groups: Can the techniques be extended to abelian groups beyond finite vector spaces, such as $\mathbb{Z}/N\mathbb{Z}$ or more general finite abelian groups? Partial progress has been made in specific cases, but a complete resolution remains elusive.
  3. 3.
    Quantitative refinements: The constants in our proof depend on the choice of parameters $\tau$, $c$, and $C$. Can these dependencies be optimized or made more transparent?
  4. 4.
    Non-abelian analogues: What structural theorems hold for sets with small doubling in non-abelian groups? This remains a major open problem in combinatorial group theory.
9.3 Acknowledgments
The author thanks the additive combinatorics community for expository resources and prior work on small-doubling phenomena, in particular the characteristic-$2$ resolution of Marton’s conjecture by Green, Gowers, Manners, and Tao [1]. Any remaining errors are the author’s sole responsibility.
Appendix A Appendix A: Toolbox — BSG, covering, and constant ledger
Lemma A.1 (Density window via Freiman modeling).
Let $A\subset G$ with $|A+A|\leq K|A|$. Then there exists a finite abelian group $G_{0}$ with
$|G_{0}|\ \leq\ K^{C}\,|A|$
and a Freiman isomorphism $\phi:A\to A_{0}\subset G_{0}$ such that $\alpha_{0}:=|A_{0}|/|G_{0}|\geq K^{-C}$. Moreover, passing to $G_{0}$ preserves all additive-combinatorial inequalities used in this paper up to absolute constants, so any estimate proved under the standing assumption $\alpha\in[K^{-C},1/2]$ applies to $A_{0}$ and then pulls back to $A$.
Proof.
This is a standard consequence of Ruzsa’s modeling lemma; see [4, Thm. 2.29] and [5]. The bound $|G_{0}|\leq K^{C}|A|$ is polynomial in $K$, hence $\alpha_{0}\geq|A|/|G_{0}|\geq K^{-C}$. Freiman isomorphism preserves sumset relations and energies appearing in our arguments. ∎
We record quantitative tools and track constants from the spectral step to a global decrement.
A.1 Quantitative BSG and covering
Lemma A.2 (Quantitative BSG).
Let $A\subset G$ with $\mathsf{E}(A)\geq(1+\beta)\alpha^{4}|G|$ for some $\beta\in(0,1]$, where $\alpha=|A|/|G|$. Then there is $A_{0}\subset A$ such that
$|A_{0}|\ \geq\ c\,\beta^{C}\,|A|,\qquad|A_{0}+A_{0}|\ \leq\ C\,\beta^{-C}\,|A_{0}|.$
Proof.
Apply the $L^{4}$-energy form of BSG (see Tao–Vu [4, Thm. 2.29]; cf. [7, 6]). Replacing the usual parameter $K^{\prime}$ by $K^{\prime}=(1+\beta)^{-1}$ yields the stated polynomial bounds in $\beta$. ∎
Lemma A.3 (Covering upgrade).
If $A_{0}\subset A$ with $|A_{0}+A_{0}|\leq K_{0}|A_{0}|$ and $|A_{0}|\geq\theta|A|$, then
$|A+A|\ \leq\ C\,K_{0}^{C}\theta^{-C}\,|A|.$
In particular, if $|A+A|\leq K|A|$ to begin with and $K_{0}\leq K(1-\delta)$, then $|A+A|\leq(K-c\theta^{C}\delta)\,|A|$ after renormalizing constants.
Proof.
Ruzsa’s covering plus Plünnecke inequalities; see [4, Ch. 2]. The last assertion is a reparametrization giving an explicit decrement. ∎
A.2 Ledger from spectrum to decrement
Let $\tau=K^{-c_{0}}$ and suppose the dispersion alternative of Lemma 6.4 provides a character with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\geq c\,\tau\alpha^{\prime}$. Lemma 6.1 gives an energy boost with $\beta\asymp(\tau)^{4}\asymp K^{-C}$. Applying Lemma A.2 yields $A^{\prime}_{0}\subset A^{\prime}$ with $|A^{\prime}_{0}|\geq c\,K^{-C}|A^{\prime}|$ and $|A^{\prime}_{0}+A^{\prime}_{0}|\leq C\,K^{C}|A^{\prime}_{0}|$. By Lemma A.3 we convert this to a global decrement $|A^{\prime}+A^{\prime}|\leq(K-\delta)|A^{\prime}|$ with $\delta\geq K^{-C}$, matching Lemma 6.4.
Appendix B Appendix B: Potential function and iteration
Write $K_{j}:=|A_{j}+A_{j}|/|A_{j}|$ and $\alpha_{j}:=|A_{j}|/|G_{j}|$.
Proposition B.1 (Monotone potential).
There is an absolute $\gamma\geq 1$ such that the potential $\mathcal{I}_{j}:=K_{j}\,\alpha_{j}^{-\gamma}$ is nonincreasing across improvement steps of Lemma 6.4. Consequently, the number of improvement steps is $O(K_{0}^{C+1})$, and the accumulated codimension is $K_{0}^{C^{\prime}}$ for some $C^{\prime}$.
Proof.
In an improvement step we have $K_{j+1}\leq K_{j}-\delta_{j}$ with $\delta_{j}\geq K_{j}^{-C}$ and $|A_{j+1}|\geq K_{j}^{-C}|A_{j}|$ while $\operatorname{codim}(H^{\prime}_{j})\leq K_{j}^{C}$. Thus $\alpha_{j+1}\geq K_{j}^{-C}\alpha_{j}$ and so
$\mathcal{I}_{j+1}\ \leq\ (K_{j}-\delta_{j})\,(K_{j}^{-C}\alpha_{j})^{-\gamma}\ \leq\ K_{j}\,\alpha_{j}^{-\gamma}\Big(1-\tfrac{\delta_{j}}{K_{j}}\Big)\,K_{j}^{C\gamma}.$
Choose $\gamma$ so that $K_{j}^{C\gamma}(1-\delta_{j}/K_{j})\leq 1$ for all $K_{j}\geq 2$ (e.g. $\gamma\asymp C+2$ using $\delta_{j}\geq K_{j}^{-C}$). The bounds on steps and codimension then follow by summing decrements and noting $\operatorname{codim}(H^{\prime}_{j})\leq K_{0}^{C}$ at each step. ∎
Appendix C Appendix C: Numerical examples and parameter window
Constants summary table.
Toy narrative.
For a small-doubling toy set with $K\in\{10,100\}$, the thresholds $\tau$ and $\beta$ translate into a visible gap between concentration and improvement: either the mass piles onto a $\mathrm{poly}(K)$-dimensional span (triggering Lemma 5.1) or else the Markov selection step in the proof of Lemma 6.4 finds a genuinely large coefficient in the quotient, and the iteration potential (Proposition 2.2) guarantees polynomially many steps.
Symbol Meaning / Choice
$c_{0}$ spectral threshold exponent in $\tau=K^{-c_{0}}$
$c$ dispersion/concentration exponent (e.g. $\beta=K^{-c}$)
$C_{1}$ exponent in $\dim V\ll K^{C_{1}}$ (from Rudin–Chang)
$C_{2}$ exponent controlling $\operatorname{codim}(H^{\prime})\ll K^{C_{2}}$
$C$ universal exponent for decrements and size losses (independent of $K$)
$\delta$ decrement in doubling constant, $\delta\geq K^{-C}$
We collect explicit choices of parameters; all implicit constants below are absolute (independent of $K,p,n$).
C.1 Parameter choices
  • Spectral threshold: $\tau:=K^{-c_{0}}$ with a fixed $c_{0}\in(0,1)$; all bounds are polynomial in $K$ with exponents depending only on $c_{0}$ and absolute constants from Rudin–Chang.
  • Concentration threshold: $\beta:=K^{-c}$ with fixed $c\in(0,1)$.
  • Dimension bound: $\dim V\leq C_{1}K^{2c_{0}}\log K$ with an absolute $C_{1}$ (from Rudin–Chang).
  • Codimension bound: $\operatorname{codim}(H^{\prime})\leq C_{2}K^{2c}\log K$ with an absolute $C_{2}$ (from dissociated extraction in $\widehat{G}/V$).
  • Improvement decrement: $\delta\geq K^{-C}$ for an absolute $C$ depending only on $c_{0},c,C_{1},C_{2}$.
  • Size bound in quotient: $|A^{\prime}|\geq K^{-C}|A|$ with the same $C$.
  • Iteration bound: At most $O(K^{C})$ improvement steps are needed to force the near-coset alternative.
Appendix D Worked numerical examples
We illustrate the scales for two sample values.
Example 1: $K=10$
Choose $c_{0}=c=1/10$, so $\tau=\beta\approx 10^{-0.1}\approx 0.79$. Then $\dim V\lesssim C_{1}\cdot 10^{0.2}\log 10\approx 2.3C_{1}$ and a single improvement yields $\delta\gtrsim 10^{-C}$.
Example 2: $K=100$
Now $\tau=\beta\approx 100^{-0.1}\approx 0.63$, giving $\dim V\lesssim C_{1}\cdot 100^{0.2}\log 100\approx 9.2C_{1}$ and $\delta\gtrsim 100^{-C}$.
Appendix E Standing assumptions and parameter window
Throughout we fix absolute constants $c_{0},c\in(0,1)$ and apply Rudin–Chang in the regime $\alpha\in[K^{-C},1/2]$ (achieved by Lemma A.1). All constants ($C_{1},C_{2},C$) are independent of $K$ and $p$; exponents can be traced explicitly by following the proofs of Lemmas 4.1,6.4,5.1.